2. Given the binary operation a*b=2a2 - b, find -2*1 (show Solution)
Answer by greenestamps(13200)



Note: Use "^" to denote exponentiation. E.g., "2a^2" for 2 a squared, instead of "2a2".

Given the definition

a*b = 2a^2-b,

if a=-2 and b=1 then

a*b = (-2)*1 = 2(-2)^2-(1) = 2(4)-1 = 8-1 = 7

This kind of statement looks strange the first couple of times you see it. But it's quite simple.

The definition "a*b = 2a^2-b" simply tells you that if you have two numbers a and b, the value of a*b is found by squaring the first number, doubling it, and subtracting the second number.